# Assessing the Alignment of FOL Closeness Metrics with Human Judgement

### v1

@@ Abstract
Ramya Keerthy Thatikonda1, Wray Buntine1,2, Ehsan Shareghi1, 1Department of Data Science & AI, Monash University, 2College of Engineering and Computer Science, VinUniversity
@@ 3 Evaluation Framework
4 Experiments
Data Preparation.
4.1 Evaluation Preparation
Perturbation.
Match $=$
Quantifier $\downarrow$
Negation $\downarrow$
AndOr $\downarrow$
OrXor $\downarrow$
Operator $\downarrow$
Predicate $=$
Variable $\downarrow$
Data (%) 100 70 99 59 32 98 22 100
LLM Samples.
Pairwise Ranking.
5 Results and Discussion
We present results from the two variations.
@@ 5 Results and Discussion
Pertb BL LE RO ME BS SP
Match $=$ 1.00 1.00 1.00 1.00 1.00 1.00
Operator Quantifier $\downarrow$ 0.96 1.00 0.96 0.96 1.00 0.99
Negation $\downarrow$ 0.69 0.73 0.93 0.85 0.97 0.37
AndOr $\downarrow$ 0.88 0.72 0.95 0.96 1.00 0.96
OrXor $\downarrow$ 0.95 0.92 0.98 0.98 1.00 0.99
Text Operator $\downarrow$ 0.20 0.62 0.53 0.42 0.89 0.44
Predicate $=$ 0.94 0.93 0.97 0.98 1.00 0.92
Variable $\downarrow$ 0.28 1.00 0.74 0.68 0.92 0.76
[…]
BL 0.96 0.94 0.96 0.96 0.97 0.93
LE 0.94 1.00 0.95 0.94 0.96 0.92
RO 0.96 0.95 0.96 0.96 0.98 0.94
ME 0.96 0.94 0.96 0.96 0.97 0.93
BS 0.97 0.96 0.98 0.97 1.00 0.95
SP 0.93 0.92 0.94 0.93 0.95 0.99

### v2

@@ Abstract
Ramya Keerthy Thatikonda1, Wray Buntine1,2, Ehsan Shareghi1 1Department of Data Science & AI, Monash University 2College of Engineering and Computer Science, VinUniversity ramya.thatikonda1@monash.edu
@@ 3 Evaluation Framework
3.1 Perturbation Evaluations
The effect of perturbations measures the sensitivity of the metrics by assessing how small changes or variations in the FOL statements impact the metric scores. Based on the ground-truth of the FOLIO dataset Han et al. (2022), we utilize nine operators to construct a formal logic framework. Perturbation variations are applied to assess the impact on metric scores. To evaluate the performance of these metrics, we first conduct a self-matching experiment on the statements and normalize the results based on the variations observed in this process. The perturbation strategies are as following:
Quantifier:
In this perturbation, we swap the quantifiers $\forall$ and $\exists$ where applicable. For example, the formula $\forall\text{x}(\text{W}(\text{x},\text{C})\rightarrow\text{A}(\text{x},\text{C}))$ becomes $\exists\text{x}(\text{W}(\text{x},\text{C})\rightarrow\text{A}(\text{x},\text{C}))$. This subtle change allows us to isolate the effect of quantifiers on the similarity metrics, demonstrating that misidentifications in quantifier use can be detected by these metrics.
Negation:
This perturbation measures the impact of negation on the metrics. We either remove the negation of predicates, if present, or add it when absent. For example, $\forall\text{x}(\neg\text{W}(\text{x},\text{C})\rightarrow\text{A}(\text{x},\text{C}))$ changes to $\forall\text{x}(\text{W}(\text{x},\text{C})\rightarrow\neg\text{A}(\text{x},\text{C}))$. This modification tests the metrics’ ability to correctly identify the placement of negations, maintaining fidelity to the underlying logic.
And/Or and Or/Xor:
This perturbation involves a simple swap of logical operators, such as (And, Or) and (Or, Xor). Given that translations by the LLMs may confuse these operators, it is important to assess how such changes are reflected in the metrics.
Operator:
This perturbation focuses on the role of operators in influencing similarity scores. All logical operators are removed, and any multiple predicates are connected by a disjunction ($\lor$) to preserve the structure. For instance, $\forall\text{x}(\neg\text{W}(\text{x},\text{C})\rightarrow\text{A}(\text{x},\text{C}))$ becomes $\text{W}(\text{x},\text{C})\lor\text{A}(\text{x},\text{C})$.
Predicate:
[…]
Variable:
[…]
3.2 Sample Evaluations
Measuring the sample correctness with respect to the ground truth allows for an assessment of alignment between different types of rankers. We randomly sampled a small set of data from FOLIO dataset and implemented a sampling process in which gpt-4o Achiam et al. (2023) was zero-shot prompted to generate three samples of FOLs for each text input. For each data point, consisting of a natural language text and its corresponding FOL label, denoted as $\{NL,FOL\}$, the $NL$ was provided to gpt-4o (see Appendix B for prompt detail) to generate three samples: $\{FOL_{1},FOL_{2},FOL_{3}\}$. In many cases, the LLM produced a correct $FOL_{1}$. To introduce randomness, we shuffled the order of the samples before passing them to the metrics. The shuffled FOL samples, $\{\hat{FOL_{1}},\hat{FOL_{2}},\hat{FOL_{3}}\}$, were then evaluated by various metrics, with a score produced for each comparison. In instances when two or all three comparisons resulted in tied scores, the ranks were adjusted to be equal. For example, if the first two FOLs were the same and the third was different, with initial ranks of [1, 2, 3], we adjusted them to [1, 1, 3].
To compare the rankings generated by these metrics, we conducted a human evaluation. We enlisted three annotators with at least a Master’s degree in CS or AI to rank the similarity between the ground truth FOL and the generated samples. The instructions provided to the experts were kept open-ended, offering only a basic overview of the logic and ranking criteria to avoid inducing bias (See Appendix C). Although the instructions suggested ranking randomly in case of a tie, we deduplicated the values and assigned the same rank to the matching FOLs, as described in the previous passage. We also used gpt-4o and o1-preview LLMs to rank the samples, allowing us to form a broader perspective on the results (see Appendix D for prompt details).
4 Experiments
Data Preparation.
Match $=$
Quantifier $\downarrow$
Negation $\downarrow$
AndOr $\downarrow$
OrXor $\downarrow$
Operator $\downarrow$
Predicate $=$
Variable $\downarrow$
Data (%) 100 70 99 59 32 98 22 100
4.1 Evaluation Preparation
Perturbation.
LLM Samples.
Pairwise Ranking.
5 Results and Discussion
We present results from the two variations.